Solve the system by substitution. Enter you answer as an ordered pair, or type none or infinite.
x=4y
2x+3y=44

Answer by sofiyac(983)

x=4y
2x+3y=44

2(4y)+3y=44
8y+3y=44
11y=44
y=44/11
y=4
x=4y=4(4)=16
(16,4)